Find all the possible values of x, y
x, y not equal to zero such that
sec2A = 4xy/(x+y)^2

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
sec^{2}A = \frac{4xy}{(x+y)^{2}}
sec^{2}A \geq 1
\frac{4xy}{(x+y)^{2}}\geq 1
4xy\geq (x+y)^{2}
(x+y)^{2}-4xy \leq 0
(x-y)^{2} \leq 0
Only solution is
x = y
x, y\in R - (0)